OpenStudy (anonymous):

Choose the correct simplification of the expression (c2)9. c18 c11 c512 c

jimthompson5910 (jim_thompson5910):

similar example \[\large (c^3)^4 = c^{3*4}\] \[\large (c^3)^4 = c^{12}\] note: this isn't the answer to your question as it's just a similar example

OpenStudy (anonymous):

what about this?Choose the correct simplification of the expression c2 • c9. c11 c7 c18 c would it be c18?

jimthompson5910 (jim_thompson5910):

no, now you're going to add the exponents (instead of multiplying them)

OpenStudy (anonymous):

so c11

jimthompson5910 (jim_thompson5910):

yep \(\large c^{11}\)

OpenStudy (anonymous):

what do you do for this? (8x2)3

jimthompson5910 (jim_thompson5910):

raise each piece to the 3rd power

jimthompson5910 (jim_thompson5910):

(8x^2)^3 = 8^3 * (x^2)^3 = ???

OpenStudy (anonymous):

512x6?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

\[\large 512x^6\]

OpenStudy (anonymous):

what do I do with this? Choose the correct simplification of the expression (x2y)2. x4y3 x6y6 x4y2 xy4

jimthompson5910 (jim_thompson5910):

square each piece

OpenStudy (anonymous):

what do you mean?

jimthompson5910 (jim_thompson5910):

square x^2 to get ____ square y to get ____

OpenStudy (anonymous):

its b

jimthompson5910 (jim_thompson5910):

(x^2)^2 = ???

OpenStudy (anonymous):

x^6

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

idk how you got 6

OpenStudy (anonymous):

multiplying 2, 2 times

jimthompson5910 (jim_thompson5910):

i see, no you just say 2*2 = 4

jimthompson5910 (jim_thompson5910):

(x^2)^2 = x^4

OpenStudy (anonymous):

xy^4?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

the 4 is for the x, not the y

OpenStudy (anonymous):

ohh so x^4y^2

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np